# Survey of Hallucination in Natural Language Generation

### v6

@@ 4.1. Statistical Metric
Given that many traditional metrics leverage the target text as the ground-truth reference (e.g., ROUGE, BLEU, etc.), Dhingra et al. (2019) build upon this idea and propose PARENT (Precision And Recall of Entailed n-grams from the Table)  [^3] Note that PARENT is a general metric like ROUGE and BLEU, not only constrained to hallucination, a metric which can also measure hallucinations using both the source and target text as references. Particularly, PARENT n-gram lexical entailment matches generated text with both the source table and target text. The F1-score that combines the precision and recall of the entailment reflects the accuracy in the table-to-text task. The source text is additionally used because it is not guaranteed that the output target text contains the complete set of information available in the input source text.
@@
Although simple and effective, one potential limitation of the lexical matching is that it can only handle the lexical information. Thus, it fails to deal with syntactic or semantic variations (Sellam et al., 2020).
@@ 8.2.4. Mitigation Methods
Rashkin et al. (2021) introduce a set of control codes and concatenate them with dialogue inputs to reduce the hallucination by forcing the model to be more aware of how the response relies on the knowledge evidence in the response generation. Some researchers have also tried to reduce hallucinated responses during generation by improving dialogue modeling. Wu et al. (2021) apply inductive attention into transformer-based dialogue models, and potentially uninformative attention links are removed with respect to a piece of pre-established structural information between the dialogue context and the provided knowledge. Instead of improving the dialogue response generation model itself, Dziri et al. (2021a) present a response refinement strategy with a token-level hallucination critic and entity-mention retriever, so that the original dialogue model is left without retraining. The former module is designed to label the hallucinated entity mentioned in the generated responses, while the retriever is trained to retrieve more faithful entities from the provided knowledge graph. RHO (ji-etal-2023-rho) is a framework that uses three mechanisms to tackle hallucinations, namely, local knowledge grounding, global knowledge grounding, and response re-ranking to tackle hallucinations in open-domain dialogues, and has been empirically shown to perform this.
@@ 13. Hallucination in Large Language Models
Scaling up model and data size for language models has shown great empirical success. The resulting Large Language Models (LLMs) not only have achieved significant performance improvements upon previous PLMs across a diverse array of NLP tasks, but have also demonstrated many emerging abilities and strong steerability after instruction tuning and Reinforcement Learning from Human Feedback (RLHF) (Zhao et al., 2023; Wei et al., 2022). However, they still exhibit the hallucination problem. Even worse, since LLMs generate highly fluent and convincing responses, their hallucinations become more difficult to identify, and more likely to have harmful consequences. The launch of ChatGPT as an LLM with a conversational user interface led to the popularity of LLMs and their wide range of real-world applications. The research on the identification and mitigation of hallucination in LLMs has intensified. LLMs are usually open-ended general-purpose systems, which differ much from task-specific models, and their architectural designs, data coverage, training methodologies, and model behaviors are also different from PLMs mentioned in previous sections. Therefore, in the following sections, we specifically discuss the hallucination problem in LLMs, by covering more recent works, introducing novel strategies for measuring and mitigating hallucination, as well as listing unsolved questions and future directions.  [^8] This section was updated in Jan 2024.
@@ 13.1. Hallucination Definition in LLMs
In the realm of LLMs, hallucination takes on a broader definition, reflecting the vastness of their training data, the breadth of their knowledge base, and their multitasking capability. Hallucination in LLMs not only signifies deviations from the source input but also extends to deviations from world knowledge. In this context, the “fact” discussed in Section 2.3 includes both the input source and the world knowledge. The hallucination degree reflects and encapsulates the model’s capacity to accurately and faithfully comprehend and represent the world. This includes a grasp of factual information, an understanding of linguistic subtleties, and an ability to navigate the intricate labyrinth of human knowledge and experience. Therefore, LLM hallucination is more oriented toward the extrinsic type involving unfaithful or nonsensical facts. Researchers explore the model’s ability to maintain the integrity of its parameterized knowledge, contextual understanding, recall relevant knowledge, and reasoning abilities (Ji et al., 2023c). This definition of hallucination in LLMs underscores the challenges and opportunities in developing these models. It highlights the need for robust mechanisms to prevent hallucination, to ensure the accuracy and reliability of the models, and to harness their potential in a responsible and beneficial manner.
@@ 13.2.1. Reference-dependent Metrics
Collecting human-verified QA samples to test the factual knowledge of LLMs is the most straightforward method. Since the training corpora of LLMs are extremely large, high frequency count knowledge and basic commonsense can be easily memorized. As a result, examining low frequency, long-tailed knowledge becomes a primary focus of these hallucination benchmarks. Testing samples for long-tail knowledge can be selected based on frequency of appearance (e.g., in pretraining corpora (Kandpal et al., 2023) or Wikitext (Du et al., 2023)), popularity measurements based on website traffic data (e.g., PopQA (Mallen et al., 2023), Head-to-Tail (Sun et al., 2023)), recency (RealTimeQA (Kasai et al., 2022)), or from specific domains (e.g., ExpertQA (Malaviya et al., 2023) and Med-HALT (Umapathi et al., 2023)). In addition to long-tail knowledge, another way to build challenging benchmarks is to gather questions that are more likely to lead to hallucinations – a methodology similar to adversarial prompting (Kumar et al., 2023; Zhu et al., 2023) and red teaming (Ganguli et al., 2022). The most representative example is TruthfulQA (Lin et al., 2022), a benchmark containing challenging questions that some humans would answer falsely due to false beliefs or misconceptions (e.g., question “If it’s cold outside, what does that tell us about global warming?” leads to GPT-3 answers “It tells us that global warming is a hoax”). Similarly, the DecodingTrust (Wang et al., 2023a) evaluated the trustworthiness of GPT models in terms of adversarial robustness, OOD knowledge, vulnerability of jail-breaking, misleading instructions, etc. Known-Unknown (Amayuelas et al., 2023) and SeflAware (Yin et al., 2023b) evaluated LLMs’ awareness of uncertainty for question without definitive answer (e.g., “If the Universe started at the Big Bang, what existed before then?”)
@@
Manual question collection and answer annotation are labor intensive, hard to scale, and hard to adapt to new domains. To address this issue, annotation-free benchmarking methods have been proposed, which can be further divided into 1) automated benchmark generation and 2) automated evidence retrieval. For the first type of methods (Muhlgay et al., 2023; Chen et al., 2023c), factual but unstructured corpora (for example, Wikipedia and reliable news articles) are transformed into a unified QA format based on information extraction and answer candidate generation. The second type of methods automatically retrieves evidence related to LLM generations from the Internet, similar to automated fact checking (Thorne and Vlachos, 2018; Guo et al., 2022). They typically adopt a chained pipeline, where search queries are first generated (by an LLM) based on the content that needs to be verified, and then a model determines whether it contains hallucinations according to the retrieved evidence (Gao et al., 2023)). This pipeline can be further enhanced by upgrading the verification model from an NLI model (Gao et al., 2023) to prompted/fine-tuned LLMs (as in AttributionScore (Yue et al., 2023) and (Varshney et al., 2023a)), or by adding additional steps to the pipeline, such as tool interaction (as in FactTool (Chern et al., 2023) and CRITIC (Gou et al., 2023a)), claim decomposition (Self-Checker (Li et al., 2023e)), additional second stage fine-grained retrieval and claim-focused summarization (as in (Chen et al., 2023a)), or chain-of-verification (Dhuliawala et al., 2023a).
@@ 13.2.2. Reference-free Metrics
Reference-free metrics aim to assess the hallucination risk of LLMs without relying on external references or ground truth. Although not as reliable as reference-based ones, these metrics are particularly useful when golden references are unavailable, expensive to obtain, or when hallucination identification needs to be performed in real time. Reference-free metrics generally fall into two categories: uncertainty-based and consistency-based hallucination detection. Uncertainty-based methods rely on the token probabilities assigned by the LLM during generation, whereas consistency-based methods evaluate the coherence of multiple completions generated by the LLM. There are also some benchmarks to measure the effectiveness of these reference-free hallucination metrics, which are also termed “hallucination detection benchmarks”, such as HaluEval (Li et al., 2023b) and HaDes (Liu et al., 2022) These benchmarks serve as meta-metrics of hallucination metrics, offering valuable resources for developing more effective methods to identify hallucinations.
@@ 13.3.1. Data-Related Methods
Recent studies (Zhou et al., 2023) showed that the knowledge within LLMs is almost entirely acquired during pretraining. Accordingly, a strong emphasis should be placed on ensuring the quality of the pretraining data. This can be performed through the collection of pretraining data from credible sources (Touvron et al., 2023a) and the minimization of defective and noisy data such as those that are unreliable or unverifiable (Zhang et al., 2023a). A representative example of such data-driven methods can be found in the development of Llama-2 (Touvron et al., 2023b), where the most factual sources within the pretraining data were up-sampled in order to reduce hallucinations. Similarly, Falcon LLM (Penedo et al., 2023) also demonstrated that their data refinement methods including rigorously filtering and deduplicating can significantly boost LLM performance. Due to the extensive pretraining datasets, some approaches have been developed to overcome the impracticalities that arise from manually selecting them. For instance, for the pretraining of GPT-3 (Brown et al., 2020), a classifier-based automatic filtering method was applied to remove low-quality documents. However, as pretraining data is being continuously scaled up, it is becoming increasingly costly for the research community to even access the entire pretraining dataset. As an alternative, small LLMs like phi-1.5 (Li et al., 2023a), which are trained on strictly controlled “textbooks-style” small-scale corpus, provide a valuable opportunity for doing rigors ablations to investigate the relationship between pretraining data and model hallucination.
@@
During the fine-tuning stage, data refinement can also be used to dampen hallucinations. As the size of fine-tuning data is significantly smaller, manual approaches can also be practically used (Zhang et al., 2023a). For example, the authors of LIMA collected a diverse dataset of 1,000 carefully curated prompts and responses which are aligned with each other (Zhou et al., 2023). Conversely, Chen et al. (2023b) point out that widely used fine-tuning datasets e.g. Alpaca’s 52k data contain low-quality instances with incorrect or irrelevant responses which are detrimental to fine-tuning, used an automated approach leveraging strong LLMs to identify and discard such low-quality data. Alongside, the authors of InstructMining (Cao et al., 2023), proposed linear rules for selecting high-quality instruction fine-tuning data and avoided the need for human or machine annotations.
@@
Training reliable reward models such that desirable and undesirable outputs are distinguished is an effective method in mitigating hallucination. These models can subsequently be incorporated into reinforcement learning pipelines. Since the effectiveness of the system is constrained by the performance of the reward model itself, emphasis has been placed on determining the most effective training methods (Lightman et al., 2023). Thus, OpenAI released a dataset, namely, PRM800K, which resulted in a SOTA reward model (Lightman et al., 2023). In the development of GPT-4 (OpenAI, 2023), two different approaches were implemented. Firstly, in tackling open-domain hallucinations, OpenAI collected real-world ChatGPT data that were flagged as non-factual by users and used it together with additional labelled comparison data to train their reward models. Secondly, for closed-domain hallucinations, OpenAI utilised GPT-4 itself to generate and subsequently utilise synthetic data for reward model training.
Information Augmentation
Hallucinations can be mitigated via the use of external knowledge, or tools. Retrieval Augmented Generation (RAG) is a framework that utilizes data to improve the quality of LLM performances through grounding the model with an external knowledge base (Martineau, 2023; Zhang et al., 2023b) which is a popular method used to mitigate hallucination (Es et al., 2023). However, although being a common approach, Zhang et al. (2023b) argue that despite improving end-task performance, RAG does not consistently mitigate the issue of hallucinations due to misalignment between user questions and stored knowledge. The LLM may, as a result, draw invalid correlations or utilize biased model knowledge to override retrieved knowledge and subsequently hallucinate. To enhance this alignment, automatic question refinement such as MixAlign (Zhang et al., 2023b) achieves this question-knowledge alignment through the use of a language model. Furthermore, it can also achieve further enhancement, if necessary, through human user clarifications. More specifically, the framework employs the LLM to map the conditions and constraints of the user question to corresponding ones in the knowledge base and in cases where this mapping process yields uncertainties or the evidence remains elusive, MixAlign generates a question seeking further clarification from the user. This approach has been shown to not only decrease model hallucinations but also improve the quality of generated responses.
Examples of leveraging external knowledge for answer refinement REFEED (Yu et al., 2023c). REFEED is an approach that first generates initial outputs, uses a retrieval model to acquire information from large document collections, and lastly refines the initial output by incorporating this retrieved information, and LLM-Augmenter, which is a plug-and-play method that uses automatic feedback. CRITIC (Gou et al., 2023b) is an example of a framework that uses external tools such as search engines and code interpreters to verify the initial output and successively amend it based on the critiques of the verification.
@@ 13.3.2. Modelling and Inference Methods
Reinforcement Learning with Human Feedback and Safety Fine-Tuning
Reinforcement learning with human feedback (RLHF) and safety fine-tuning are one of the most dominant methods utilized to mitigate hallucination. One notable example is the resultant model of GPT-3 having undergone fine-tuning and further fine tuning using supervised learning and RLHF respectively, coined as InstructGPT, which roughly halved the hallucination rate from 41% to 21% in closed domain tasks such as summarization and closed domain QA (Ouyang et al., 2022). In another well-known study, the authors of Llama-2-Chat (Touvron et al., 2023b) observed a compelling finding, which is that after various RLHF iterations, despite the rise in temperature of the LLM, for prompts based on factual information, the model learns to consistently provide the same factual response whilst retaining its capability in generating diversity when fed creativity-based prompts. Furthermore, the authors utilize RLHF with the aim to improve the helpfulness (defined as how well the LLM responses fulfill users’ requests and provide requested information) and safety of the LLM (Touvron et al., 2023b), which can decrease hallucination. Moreover, supervised safety fine-tuning was also carried out to improve model safety. It should be noted that the criterion for safety in this study not only involves whether harmful responses were produced, but also unhelpful ones (Touvron et al., 2023b). As part of RLHF, the reward models were initialized from pretrained chat model checkpoints to ensure that both the chat model and reward model benefit from knowledge acquired form pretraining thereby avoiding an information mismatch which could give rise to hallucinations (Touvron et al., 2023b).
@@
The modular architecture of LLMs can also be leveraged to combat hallucination. For example, researchers from MIT and Microsoft propose an approach, namely, Decoding by Contrasting Layers (DoLa), which emphasises knowledge from higher layers of an LLM whilst downplaying lower or intermediate layer knowledge (Chuang et al., 2023). Since factual knowledge is progressively injected by higher layers, the probability of the correct next word increases further up the architecture. Therefore, by exploiting the modularity of knowledge encoding in LLMs, the authors propose to amplify factual knowledge of LLMs via obtaining the probability of the next word from the difference between logits produced by a higher and lower layer. By making this contrast, the prominence of the correct next word’s probability is increased. Results from the study demonstrate that DoLa consistently improves truthfulness which consequently reduces hallucination.
Similarly, Context-Aware Decoding (CAD) (Shi et al., 2023), also exploits logits and uses a contrastive approach to mitigate hallucination. However, instead of contrasting between layers, CAD amplifies the difference between the output probabilities when a model is used with context and without. This approach effectively suppresses prior knowledge when contextual information is provided.
Model Editing
Subtracting the parameters of a “negative expert” fine-tuned on negative samples from a pre-trained LLM can also serve to reduce hallucination (Daheim et al., 2023). However, given that some parameters are more responsible than others in causing hallucination, Daheim et al. (Daheim et al., 2023) proposed an improved method coined as Elastic Weight Removal (EWR), which uses an approximation of the Fisher Information matrix, to weigh their individual performance, and computes a vector of differences between pre-trained and hallucination (anti-)expert models (trained with synthetic data) based on the weighting. Similarly, parameter interpolations with an abstractiveness expert model were also performed to prevent an increase in extractives, which is a potential side-effect of discouraging hallucinations. This steers the model towards being able to re-phrase ground-truth knowledge and drawing new valid conclusions rather than directly copying sections from documents with a limited understanding of their meaning.
Based on the notion that models “know” more than they “say”, the researchers at Harvard (Li et al., 2023d) studied the difference between generation accuracy determined by a model’s output and the probe accuracy (i.e. sentence classification using a model’s intermediate activations as input to a classifier). The difference observed was 40%, which shows a major gap between the information present at intermediate layers and the output. Thus, to narrow this gap, the authors propose Inference-Time intervention, a computationally inexpensive and data-efficient technique that requires only a few hundred samples. The approach involves first using supervised learning to identify attention heads with high linear probing accuracy for truthfulness, in other words, identifying latent vectors related to factual outputs using supervised learning. And secondly, during LLM inference, shift model activations along these truth-correlated directions. The same intervention is repeated autoregressively until the entire answer is generated. ITI was shown to improve truthfulness from 32.5% to 65.1% on Alpaca, an instruction fine-tuned Llama, which suggests the effectiveness of this approach in mitigating hallucinations.
@@
Chain-of-Thought can be leveraged to reduce hallucinations. For example, simply adding the prompt, “Let’s think step by step”, encourages LLMs to reason prior to arriving to an answer, which improves model performance (Kojima et al., 2023). This is further evidenced in a related study where error rates were significantly reduced through using the same prompting strategy (Zhang et al., 2023c).
At inference, reasoning-based methods such as Chain-of-Verification (COVE) (Dhuliawala et al., 2023b) and Chain of Natural Language Inference (CoNLI) (Lei et al., 2023) have been developed to reduce hallucination. The COVE method consists of four steps where an LLM first drafts an initial response, plans verification questions to fact check this, answers the questions independently to ensure the answers are not biased by other responses, and finally generates its final verified response. The framework, CoNLI, involves three steps, that is, to firstly select sentences (considered in this case as a set of hypotheses), secondly detect hallucinations with sentence and entity level detectors through prompting the LLM to solve a sequence of inference problems, and lastly to use a mitigation agent which uses detection reasonings as mitigation instructions to produce a refined response.
Self-reflection and self-collaboration is another approach that can be used to combat hallucination (Ji et al., 2023c; Shinn et al., 2023; Wang et al., 2023b). Ji et al. (2023c) is an iterative feedback loop process that involves generating relevant background knowledge for a given question followed by a factuality evaluation. When discrepancies are detected, the model is made to self-correct, which leverages its inherent capacity to reflect and thereby refine the knowledge. This iteration is continued until a satisfactory level of factuality is attained. Furthermore, during the answering stage, a similar generation-score-refine strategy is used along with an entailment evaluation to determine whether the answer properly addresses the question. If the answer fails to do so, the process returns back to the initial stage and repeats the cycle. This dynamic interaction between the system and its knowledge enhances the model to provide accurate, reliable, as well as factually grounded responses. A Solo Performance Prompting (SPP) (Wang et al., 2023b) is proposed in light of LLMs’ tendency to struggle with tasks that require intensive domain knowledge and complex reasoning. Through prompting the LLM to engage in multi-turn self-collaboration with multiple personas, SPP transforms a sole LLM into an intelligent agent that works with multiple minds such that their individual strengths and knowledge are combined in order to boost overall problem-solving and performance in such tasks.
@@
Ensembling multiple models to mitigate hallucination is another way. The “Society of Minds” approach can be used as an alternative to methods that are data-driven, modeling-driven, or at inference. One example of such (multiagentdebate2023), where multiple LLMs individually produce a response and then jointly debate each other on the responses and reasoning, such that a single common answer is reached, has been demonstrated to improve the factuality of generated content and reduce hallucination. This multi-agent interaction more specifically involves multiple instances of an LLM to first generate individual candidate answers to a given query. Subsequently, each individual model instance reads and critiques the responses of all other models and uses this content to refine its own answer; a step which is reiterated several times. This entire process steers the models towards constructing answers that are consistent with their own internal criticism and also sensible provided the response from other agents.
Post-processing
PURR (Petite Unsupervised Research and Revision) (Chen et al., 2023e), requires LLMs to introduce corruptions/noises into data and subsequently fine-tune an editor to denoise these corruptions. Specifically, PURR, a fusion-in-decoder T5 model, is trained to denoise corruptions when simultaneously given relevant evidence. Given an ungrounded output generated by LLMs, PURR generates questions to search for relevant evidence, which is subsequently used to produce an edit for the output.

### v7

@@ 4.1. Statistical Metric
Given that many traditional metrics leverage the target text as the ground-truth reference (e.g., ROUGE, BLEU, etc.), Dhingra et al. (2019) build upon this idea and propose PARENT (Precision And Recall of Entailed n-grams from the Table)  [^3] Note that PARENT is a general metric like ROUGE and BLEU, not only constrained to hallucination, a metric which can also measure hallucinations using both the source and target text as references. Particularly, PARENT n-gram lexical entailment matches generated text with both the source table and target text. The F1-score that combines the precision and recall of the entailment reflects the accuracy of the table-to-text task. The source text is additionally used because it is not guaranteed that the output target text contains the complete set of information available in the input source text.
@@
Although simple and effective, one potential limitation of lexical matching is that it can only handle lexical information. Thus, it fails to deal with syntactic or semantic variations (Sellam et al., 2020).
@@ 8.2.4. Mitigation Methods
Rashkin et al. (2021) introduce a set of control codes and concatenate them with dialogue inputs to reduce the hallucination by forcing the model to be more aware of how the response relies on the knowledge evidence in the response generation. Some researchers have also tried to reduce hallucinated responses during generation by improving dialogue modeling. Wu et al. (2021) apply inductive attention into transformer-based dialogue models, and potentially uninformative attention links are removed with respect to a piece of pre-established structural information between the dialogue context and the provided knowledge. Instead of improving the dialogue response generation model itself, Dziri et al. (2021a) present a response refinement strategy with a token-level hallucination critic and entity-mention retriever, so that the original dialogue model is left without retraining. The former module is designed to label the hallucinated entity mentioned in the generated responses, while the retriever is trained to retrieve more faithful entities from the provided knowledge graph. RHO (Ji et al., 2023a) is a framework that uses three mechanisms to tackle hallucinations, namely, local knowledge grounding, global knowledge grounding, and response re-ranking to tackle hallucinations in open-domain dialogues, and has been empirically shown to perform this.
@@ 13. Hallucination in Large Language Models
Scaling up models and data sizes for language models has shown great empirical success. The resulting Large Language Models (LLMs) not only have achieved significant performance improvements upon previous PLMs across a diverse array of NLP tasks, but have also demonstrated many emerging abilities and strong steerability after instruction tuning and Reinforcement Learning from Human Feedback (RLHF) (Zhao et al., 2023; Wei et al., 2022). However, they still exhibit the hallucination problem. Even worse, since LLMs generate highly fluent and convincing responses, their hallucinations become more difficult to identify, and more likely to have harmful consequences. The launch of ChatGPT as an LLM with a conversational user interface led to the popularity of LLMs and their wide range of real-world applications. The research on the identification and mitigation of hallucination in LLMs has intensified. LLMs are usually open-ended general-purpose systems, which differ much from task-specific models, and their architectural designs, data coverage, training methodologies, and model behaviors are also different from PLMs mentioned in previous sections. Therefore, in the following sections, we specifically discuss the hallucination problem in LLMs, by covering more recent works, introducing novel strategies for measuring and mitigating hallucination, as well as listing unsolved questions and future directions.  [^8] This section was updated in Jan 2024.
@@ 13.1. Hallucination Definition in LLMs
In the realm of LLMs, hallucination takes on a broader definition than before because of the vastness of training data, the breadth of knowledge base, and multitasking capability. Unlike earlier NLG models that depend heavily on the input source, LLMs leverage their vast internal parametric knowledge. Consequently, hallucination in LLMs not only signifies deviations from the source input but also extends to deviations from training data, marking it more oriented towards the Extrinsic type. In other words, the reference discussed in Section 2.3 includes both the input source and the training data.
Existing works often assume that training data is consistently factual and treat it as a stand-in for “fact” or “truth”. This assumption allows for a practical simplification, with the degree of hallucination reflecting the model’s ability to accurately and faithfully comprehend and represent the world. However, this assumption is not always valid, given that training datasets may contain non-factual and contradictory information. Thus, the concept of hallucination should be distinguished from that of factuality.
@@ 13.2.1. Reference-dependent Metrics
Collecting human-verified QA samples to test the factual knowledge of LLMs is the most straightforward method. Since the training corpora of LLMs are extremely large, high frequency count knowledge and basic commonsense can be easily memorized. As a result, examining low frequency, long-tailed knowledge becomes a primary focus of these hallucination benchmarks. Testing samples for long-tail knowledge can be selected based on frequency of appearance (e.g., in pretraining corpora (Kandpal et al., 2023) or Wikitext (Du et al., 2023b)), popularity measurements based on website traffic data (e.g., PopQA (Mallen et al., 2023), Head-to-Tail (Sun et al., 2023)), recency (RealTimeQA (Kasai et al., 2022)), or from specific domains (e.g., ExpertQA (Malaviya et al., 2023) and Med-HALT (Umapathi et al., 2023)). In addition to long-tail knowledge, another way to build challenging benchmarks is to gather questions that are more likely to lead to hallucinations – a methodology similar to adversarial prompting (Kumar et al., 2023; Zhu et al., 2023) and red teaming (Ganguli et al., 2022). The most representative example is TruthfulQA (Lin et al., 2022), a benchmark containing challenging questions that some humans would answer falsely due to false beliefs or misconceptions (e.g., question “If it’s cold outside, what does that tell us about global warming?” leads to GPT-3 answers “It tells us that global warming is a hoax”). Similarly, the DecodingTrust (Wang et al., 2023a) evaluated the trustworthiness of GPT models in terms of adversarial robustness, OOD knowledge, vulnerability of jail-breaking, misleading instructions, etc. Known-Unknown (Amayuelas et al., 2023; Azaria and Mitchell, 2023) and SeflAware (Yin et al., 2023b) evaluated LLMs’ awareness of uncertainty for question without definitive answer (e.g., “If the Universe started at the Big Bang, what existed before then?”)
@@
Manual question collection and answer annotation are labor intensive, hard to scale, and hard to adapt to new domains. To address this issue, annotation-free benchmarking methods have been proposed, which can be further divided into 1) automated benchmark generation and 2) automated evidence retrieval. For the first type of methods (Muhlgay et al., 2023; Chen et al., 2023c), factual but unstructured corpora (for example, Wikipedia and reliable news articles) are transformed into a unified QA format based on information extraction and answer candidate generation. The second type of methods automatically retrieves evidence related to LLM generations from the Internet, similar to automated fact checking (Thorne and Vlachos, 2018; Guo et al., 2022). They typically adopt a chained pipeline, where search queries are first generated (by an LLM) based on the content that needs to be verified, and then a model determines whether it contains hallucinations according to the retrieved evidence (Gao et al., 2023)). This pipeline can be further enhanced by upgrading the verification model from an NLI model (Gao et al., 2023) to prompted/fine-tuned LLMs (as in AttributionScore (Yue et al., 2023) and (Varshney et al., 2023a)), or by adding additional steps to the pipeline, such as tool interaction (as in FactTool (Chern et al., 2023) and CRITIC (Gou et al., 2023)), claim decomposition (Self-Checker (Li et al., 2023d)), additional second stage fine-grained retrieval and claim-focused summarization (as in (Chen et al., 2023a)), or chain-of-verification (Dhuliawala et al., 2023).
@@ 13.2.2. Reference-free Metrics
Reference-free metrics aim to assess the hallucination risk of LLMs without relying on external references or ground truth. Although not as reliable as reference-based ones, these metrics are particularly useful when golden references are unavailable, expensive to obtain, or when hallucination identification needs to be performed in real-time. Reference-free metrics generally fall into two categories: uncertainty-based and consistency-based hallucination detection. Uncertainty-based methods rely on the token probabilities assigned by the LLM during generation, whereas consistency-based methods evaluate the coherence of multiple completions generated by the LLM. There are also some benchmarks to measure the effectiveness of these reference-free hallucination metrics, which are also termed “hallucination detection benchmarks”, such as HaluEval (Li et al., 2023b) and HaDes (Liu et al., 2022) These benchmarks serve as meta-metrics of hallucination metrics, offering valuable resources for developing more effective methods to identify hallucinations.
@@ 13.3.1. Data-Related Methods
Recent studies (Zhou et al., 2024) show that the knowledge within LLMs is almost entirely acquired during pretraining. Thus, a strong emphasis should be placed on ensuring the quality of the pretraining data. This can be performed through the collection of pretraining data from credible sources (Touvron et al., 2023a) and the minimization of defective and noisy data such as those that are unreliable or unverifiable (Zhang et al., 2023a). A representative example of such data-driven methods can be found in the development of Llama-2 (Touvron et al., 2023b), where the most factual sources within the pretraining data were up-sampled in order to reduce hallucinations. Similarly, Falcon LLM (Penedo et al., 2023) also demonstrated that their data refinement methods including rigorously filtering and deduplicating can significantly boost LLM performance. Due to the extensive pretraining datasets, some approaches have been developed to overcome the impracticalities that arise from manually selecting them. For instance, for the pretraining of GPT-3 (Brown et al., 2020), a classifier-based automatic filtering method was applied to remove low-quality documents. However, as pretraining data is being continuously scaled up, it is becoming increasingly costly for the research community to even access the entire pretraining dataset. As an alternative, small LLMs like phi-1.5 (Li et al., 2023a), which are trained on strictly controlled “textbooks-style” small-scale corpus, provide a valuable opportunity for doing rigors ablations to investigate the relationship between pretraining data and model hallucination.
@@
During the fine-tuning stage, data refinement can also be used to dampen hallucinations. As the size of fine-tuning data is significantly smaller, manual approaches can also be practically used (Zhang et al., 2023a). For example, the authors of LIMA collected a diverse dataset of 1,000 carefully curated prompts and responses which are aligned with each other (Zhou et al., 2024). Conversely, Chen et al. (2023b) point out that widely used fine-tuning datasets e.g. Alpaca’s 52k data contain low-quality instances with incorrect or irrelevant responses which are detrimental to fine-tuning, used an automated approach leveraging strong LLMs to identify and discard such low-quality data. Alongside, the authors of InstructMining (Cao et al., 2023), proposed linear rules for selecting high-quality instruction fine-tuning data and avoided the need for human or machine annotations.
@@
Training reliable reward models such that desirable and undesirable outputs are distinguished is an effective method for mitigating hallucination. These models can subsequently be incorporated into reinforcement learning pipelines. Since the effectiveness of the system is constrained by the performance of the reward model itself, emphasis has been placed on determining the most effective training methods (Lightman et al., 2023). Thus, OpenAI released a dataset, namely, PRM800K, which resulted in a SOTA reward model (Lightman et al., 2023). In the development of GPT-4 (OpenAI, 2023), two different approaches were implemented. Firstly, in tackling open-domain hallucinations, OpenAI collected real-world ChatGPT data that were flagged as non-factual by users and used it together with additional labeled comparison data to train their reward models. Secondly, for closed-domain hallucinations, OpenAI utilized GPT-4 itself to generate and subsequently utilize synthetic data for reward model training.
Retrieval Augmented Generation (RAG)
RAG is a technique that leverages external sources to enhance the reliability of generative models (Martineau, 2023; Zhang et al., 2023b), which is a popular method used to mitigate hallucination (Es et al., 2023). For example, Gemini incorporates the search engine to provide external references during generation. However, RAG faces several challenges (Chen et al., 2023c): In the retrieval stage, the model is required to discern the noise, irrelevant or fake information. In the augmentation stage, integrating heterogeneous, independent, and even conflicting information presents challenges. Additionally, the model needs to reject generation when insufficient information is given. Recent works make efforts to address these challenges. For example, MixAlign Zhang et al. (2023b) enhances the alignment between user questions and stored knowledge to improve retrieval and reduce hallucination.
@@ 13.3.2. Modelling and Inference Methods
Safety Fine-Tuning and Reinforcement Learning with Human Feedback (RLHF)
Safety fine-tuning is a prevalent method to mitigate risks associated with LLMs, such as hallucinations. A prominent technique in this domain is Reinforcement Learning with Human Feedback (RLHF). The application of RLHF in the fine-tuning process of GPT-3 leads to the development of InstructGPT, which significantly reduces the hallucination rate from 41% to 21% in closed domain tasks (Ouyang et al., 2022). Similarly, Llama2 (Touvron et al., 2023b) demonstrates that post-RLHF, the model consistently generated factual responses to prompts on factual information. The study also highlighted the use of RLHF in enhancing the overall safety of LLMs, including a reduction in hallucinations (Touvron et al., 2023b). Despite the advantages, Safety Fine-Tuning and RLHF can incur ‘alignment tax’ and ‘catastrophic forgetting’ where the models lose diverse previously acquired abilities after alignment (Lin et al., 2023). To counterbalance these issues, Lin et al. (2023) proposes Adaptive Model Averaging (AMA) optimizing the averaging ratios of different model layers to maximize alignment reward.
Model Editing
Editing Models such as parameter adapting can also serve to reduce hallucination. Given that some parameters are more responsible for causing hallucinations, EWR (Elastic Weight Removal) (Daheim et al., 2023) weighs their individual importance via Fisher Information matrix and performs parameter interpolation to remove the undesired behaviors.
@@
Without editing architecture or additional fine-tuning, novel decoding strategies are proposed to reduce hallucinations. Inspired by the finding that factual knowledge is localized to particular layers, DoLa (Decoding by Contrasting Layers) (Chuang et al., 2023) utilizes the difference in logits from the later layers versus earlier layers to amplify the factual knowledge probability. Similarly, CAD (Context-Aware Decoding) (Shi et al., 2023) exploits the difference in logits with and without context. In addition, Li et al. (2024) observe the gap between surface generation and internal knowledge and propose ITI (Inference-Time Intervention) to reduce this gap. It first identifies attention heads with high linear probing accuracy for truthfulness and then shifts activations along these truth-correlated directions during inference.
@@
Chain-of-Thought and its variants encourage LLMs to reason prior to arriving at a final reply, which improves performance and reduces hallucinations by leveraging the model’s internal knowledge (Kojima et al., 2023; Dhuliawala et al., 2023; Lei et al., 2023; Ji et al., 2023b; Shinn et al., 2023; Wang et al., 2023b). For example, Chain-of-Verification (COVE) (Dhuliawala et al., 2023) consists of four steps where an LLM first drafts an initial response, plans verification questions to fact check it, answers the questions independently to ensure the answers are not biased by other responses, and finally generates the verified response. Chain of Natural Language Inference (CoNLI) (Lei et al., 2023) detect hallucinations in the initial response by prompting the LLM to conduct sentence/entity-level NLI, and then refine the response according to the detection result. Ji et al. (2023b) propose iterative self-reflection loops that involve generating relevant background knowledge for a given question followed by a factuality evaluation and self-correction. During answering, a similar generation-score-refine strategy is used for its entailment with the question. Solo Performance Prompting (SPP) (Wang et al., 2023b) prompts a single LLM to engage in multi-turn self-collaboration with multiple personas, such that their individual strengths and knowledge are combined to boost the problem-solving and performance in tasks.
Post-processing
Petite Unsupervised Research and Revision (PURR) (Chen et al., 2023e) fine-tune an editor to denoise the corruptions and use the editor to correct the hallucinations in the initial outputs. Other works like REFEED (Yu et al., 2023c) leverage retrieved documents for post-correction. Similarly, CRITIC (Gou et al., 2023) uses external tools such as search engines and code interpreters to verify and post-correct the initial output.
@@
Ensembling multiple models to mitigate hallucination is another way. For example, in (Du et al., 2023a), multiple LLMs individually produce responses and then jointly debate their responses and reasoning, ultimately reaching a single common answer. This approach has been demonstrated to improve the factuality of generated content.